Construct the 2 by 2 table for CMZ test the gold standard when a cutoff of >41 units is used for classification ( equal or greater).

A new, rapid blood test [CMZ test] was developed to test febrile patients for viral infection. You are asked to assist with the validation. The CMZ assay was run on 789 febrile people who arrived at the local public hospital seeking medical care. Samples from patients were divided into Opox infected and not infected using a combination of PCR and virus isolation methods, which can be considered a gold standard. eet res(ity units on bto 789 people. Gold standard No Infection Qpox Infection CMZ Test result (Units) 0-20 21-40 41-60 61-80 81-100 Total number of patients No. patients No. patients 359 186 75 0 32 41 89 167 0 622 1. Construct the 2 by 2 table for CMZ test versus the gold standard when cutoff of 241 units is used for classification (equal or greater is positive)
Answer #1

There are 6steps for creating 2 by 2table for CMZ test:

- create a observation frequency

- calculate the column and row total

- calculate the expected frequency

- calculate the sum difference

- compare the values

In gold standard number of not infected patients 359 +186+75+2+-0= 622

Total number of Qpox infection Patient 0+5+32+41+89 = 167

0-20=359

21-40= 191

41-60 = 107

61-80 = 43

81- 100= 89

Total = 789

They used cutoff>41 units for the combination of CMZ test and gold standard..it was normal after cutoff >41units..

It is equal